A physicist at a fireworks display times the lag between seeing an explosion and hearing its sound, and finds it to be 0.300 s. (Enter your answers to at least four decimal places.) (a) How far away (in m) is the explosion if air temperature is 19.0°C and if you neglect the time taken for light to reach the physicist? 102.5998 (b) Calculate the distance to the explosion (in m) taking the speed of light into account. Note that this distance is negligibly greater. 0.000118 ×m
Answer #1

(a)velocity of sound at 19oC = 341.4 m/s

t=0.3 s

distance= vxt

d=341.4 (0.3)

d=102.42m

(b) time taken by light to recah the physicst= d/c=d/(3x108)

time taken by sound to reach the physicst= d/v= d/(341.4)

so d/(341.4)-d/(3x108)=0.3

d= 102.4201166m